The lengths of the sides of a triangle are 12,15,and 20.If the length of the shortest side of a similar triangle is 3, find the length of the longest side of that triangle.
Answer by edjones(8007)

Let x=length of the longest side of the smaller triangle.
x/20=3/12
x=60/12
x=5
.
Ed